What is y=x^(2)+8x-1 in vertex form?
Answer by lwsshak3(11628)

What is y=x^(2)+8x-1 in vertex form?
**
y=x^(2)+8x-1
complete the square
y=(x^2+8x+16)-1-16
y=(x+4)^2-17
This is an equation of a parabola that opens upwards with vertex at (-4,-17)
Its standard form of equation:y= (x+h)^2+k, (h,k) =(x,y) coordinates of the vertex.